OpenStudy (anonymous):

A gambler started playing a game.in the first round he trippled his amount and gave p rs to his friend. In second round he doubled his amount what he have and gave 3p to his friend and in third round he quadrapauled the amount and gave 5p to his friend.at the end he was left with no money.if he gave total 360 rs to his friend then how much money with he did start palying?

OpenStudy (anonymous):

plz help me with this que..

OpenStudy (anonymous):

first calculate the value of p given p+3p+5p=360 and calculate backward...

OpenStudy (anonymous):

that is equal to 40 right?..bt how do i form the equation for the amount he started for plaaying?

OpenStudy (anonymous):

yup 40 is correct

OpenStudy (anonymous):

so calulate backward he finally gave 5p=200 that means after second round he must have with him 200/4=50 and so on...

OpenStudy (debbieg):

You need 2 equations, one is as @matricked said above, p+3p+5p=360. Then use another variable, say A, for the amount he started with, and work with that to build the other equation, which will model his process of tripling then giving away, doubling then giving away, and so on. So first he tripled his amount, that's 3A Then he gave away p, so you get 3A-p Next he doubled what he has at THAT point, so he has 2(3A-P) Can you work it out from there?

OpenStudy (anonymous):

so according to it it should be like in first round=3A-p,2nd round=6A-9p and in 3rd round=24A-13p is it right?..now what?

OpenStudy (anonymous):

hey i got it...u two are amazing..thk u both debbie and matrixed..